# CodeIF-Bench: Evaluating Instruction-Following Capabilities of Large Language Models in Interactive Code Generation

### v3

@@
Abstract.
Large Language Models (LLMs) have demonstrated exceptional performance in code generation tasks and have become indispensable programming assistants for developers. However, existing code generation benchmarks primarily assess the functional correctness of code generated by LLMs in single-turn interactions. They offer limited insight into LLMs’ abilities to generate code that strictly follows users’ instructions in multi-turn interaction scenarios. In this paper, we introduce CodeIF-Bench, a benchmark for evaluating the instruction-following capabilities of LLMs in interactive code generation. Specifically, CodeIF-Bench incorporates nine types of verifiable instructions aligned with the real-world software development requirements, which can be independently and objectively validated through specified test cases, facilitating the evaluation of instruction-following capability in multi-turn interactions. In both Static Conversation and Dynamic Conversation settings, we evaluate the performance of 7 state-of-the-art LLMs and summarize the important factors influencing the instruction-following ability of LLMs in multi-turn interactions, as well as potential directions for improvement.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineeringccs: Computing methodologies Artificial intelligencecopyright: none
1. Introduction
In recent years, the remarkable advancements in Large Language Models (LLMs) for code generation (Zhu et al., 2024; Roziere et al., 2023; Luo et al., 2023) have significantly enhanced developers’ coding efficiency through an interactive and conversational paradigm (Github, 2021; Anysphere, 2023). In practical software development environments, as shown in Figure 1, the collaboration between developers and LLMs typically necessitates an interactive code generation process. Throughout this interactive workflow, developers frequently enhance and refine their initial instructions by adding supplementary details to better align with their requirements. Specifically, through systematic code execution and rigorous review processes, developers often identify and provide follow-up instructions to clarify requirements and address issues in LLM-generated code. These issues may include implementing new features (Jimenez et al., 2023), fixing bugs (Tambon et al., 2024), resolving potential security vulnerabilities (Perry et al., 2023), etc. Specifically, any discrepancies or misunderstandings in LLM following instructions can significantly increase interaction time and cost, as developers may need to spend additional time clarifying and fixing errors. Besides, such misunderstandings might even trigger uncontrollable and unexpected issues, such as bugs (Tambon et al., 2024), system failures (Dunne et al., 2024), or even security vulnerabilities (Perry et al., 2023), which can have severe implications for the software’s performance and developers’ trust. Consequently, a dedicated benchmark to evaluate LLMs’ Instruction-Following (IF) capabilities in interactive code generation is essential.
@@ 1. Introduction
Figure 1. An example of interactive code generation, where the developer provides follow-up instructions to clarify the requirement and address issues in the generated code.
The key challenges for a benchmark to evaluate the instruction-following capabilities of LLMs in interactive code generation tasks include:
  • Challenge 1: The benchmark data should support the multi-turn interaction with LLM in code generation.
  • Challenge 2: The inherent complexity of natural language makes evaluation difficult.
  • Challenge 3: The benchmark data should align with real-world development scenarios.
@@
For Challenge 1, existing popular code generation benchmarks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Li et al., 2024a) primarily evaluate whether LLMs can generate functionally correct code given specific programming tasks in a single turn, and it is difficult to directly evaluate LLMs’ instruction-following capability in multi-round interaction. For Challenge 2, for instance, subjective instructions such as “Please make the comments clearer” may result in varying interpretations when assessing whether the model follows the given requirment. Some instruction-following benchmarks in the NLP domain, including for multi-turn (He et al., 2024; Zheng et al., 2023; Bai et al., 2024), most use the LLM-as-Judge method for evaluation, which may lead to bias (Li et al., 2025; Chen et al., 2024a). To circumvent this challenge, some work, such as (Zhou et al., 2023), focuses on evaluating the ability of LLMs to follow “verifiable instructions” which are defined as instructions that can be objectively verified for compliance, e.g., “the response should be more than 300 words”. However, it does not address the Challenge 3: these benchmarks do not align with real-world software development scenarios and fail to evaluate LLM’s ability to follow instructions in realistic interactive coding tasks.
@@
To address the above challenges, we propose a new benchmark named CodeIF-Bench, which aims to evaluate LLMs’ instruction-following capability in interactive code generation. We first gather initial instructions from real-world projects (Li et al., 2024a) and crowd-sourced problems (Austin et al., 2021), enabling evaluation for both standalone function and repository-level code generation. Then we use a combination of LLM and manual annotation to extract nine primary verifiable instructions strategies of real-world user requirements to guide LLM to generate verifiable instructions, which can be independently and objectively verified using their respective test cases, such as initial instructions. We guide LLM to generate verifiable instructions based on these strategies and initial instructions, resulting in each data point in CodeIF-Bench containing at least 8 verifiable instructions (8 for function-level and 10 for repository-level) to construct a conversation. In addition, we design Static Conversation (predefined instruction sequences) and Dynamic Conversation (providing feedback information), and evaluation metrics based on test cases to evaluate the performance of LLM in code interaction scenarios.
We evaluate the IF capabilities of 7 widely-used and advanced LLMs (i.e. GPT-4o (OpenAI et al., 2024), Claude-3.5-Sonnet (claude, 2023), DeepSeek-V3 (Zhu et al., 2024), Qwen-Coder-Turbo, Qwen2.5-Coder-{7B, 14B, 32B} (Hui et al., 2024)) in interactive code generation tasks using CodeIF-Bench. Experimental results show the key factors, such as long context and the feedback utilization capability, influencing LLMs’ instruction-following capabilities in interactive code generation tasks and potential methods for improvement. In summary, our contributions are as follows:
@@
  • We perform a thorough evaluation of 7 prominent LLMs on CodeIF-Bench, analyzing the key factors influencing their IF capabilities in interactive code generation.
2. Related Work
Code Generation Benchmarks. Recent years have seen remarkable progress in LLM-based code generation, with models such as DeepSeek-Coder (Zhu et al., 2024), WizardCoder (Luo et al., 2023), and Qwen-Coder (Hui et al., 2024) demonstrating exceptional performance. This progress has facilitated the development of interactive coding tools (Anysphere, 2023; Github, 2021) that enhance programmer productivity through effective human-LLM collaboration. This development has driven the development of benchmarks in code generation. These include benchmarks for evaluating LLMs’ performance on various code generation scenarios, including stand-alone function generation (e.g., HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021)), class-level code generation (e.g., ClassEval (Du et al., 2023)), and repository-level code generation (e.g., DevEval (Li et al., 2024a) and CoderEval (Yu et al., 2024), which leverage real-world open-source repositories). While these benchmarks primarily assess whether LLMs can generate functionally correct code in single-turn interactions, they are insufficient for evaluating LLMs’ interactive capabilities, particularly in multi-turn interaction scenarios.
Instruction Following Benchmarks. Current benchmarks for evaluating the instruction-following capabilities of LLMs (Zhou et al., 2023; Kwan et al., 2024; Zheng et al., 2023; Li et al., 2024b) primarily focus on natural language QA tasks in the NLP domain. However, there is a significant disparity exists between code generation and QA tasks. Some of the existing benchmarks utilize LLM-as-judge to assess LLMs’ instruction following performance (Wei et al., 2024; Huang et al., 2024; Chen et al., 2024b; Zheng et al., 2023), which typically rely on elaborate prompts and can introduce inherent bias. CanItEdit (Cassano et al., 2023) is a code editing benchmark, but its manually crafted programs lack project-specific dependencies, misaligning with real-world development. Additionally, its single-round dialogue design cannot assess LLMs’ performance in complex, multi-turn interactions. CodeIF (Yan et al., 2025) benchmarks instruction-following in code generation but oversimplifies real-world complexity. Its single-turn setting also fails to assess LLMs’ multi-turn interaction capabilities.
3. CodeIF-Bench
@@
Figure 2. CodeIF-Bench construction pipeline. The top part illustrates the verifiable instruction strategy extraction process, and the bottom part presents the data collection procedure.
3.1. Task Definition
@@
[…]
Static Conversation: In this scenario, we simulate a linear progression of requirements where instruction sequences are predetermined. The LLM generates code for each instruction in sequence without iterative refinement, testing its ability to maintain consistency in multi-turn interactions without external feedback. In an N-Round conversation, we initialize the process at $N=1$ with the original programming instruction ($I_{1}$) from the benchmark. Subsequent instructions are then added sequentially from a predefined non-conflicting instruction sequence as $N$ increases. In particular, for non-standalone tasks, which contain context-aware dependencies in repositories, $I_{1}$ consists of both the original task instruction and the golden contexts. Specifically, in $i$-th round, the input is formed as $\textit{Input}=(I_{1},A_{1},I_{2},A_{2},...,I_{i})$, LLM’s answer in this round is $\textit{A}_{i}$ evaluated with the test cases $(T_{1},T_{2},..T_{i})$.
Dynamic Conversation: In real-world coding scenario, programmers typically review LLM-generated outputs to provide feedback. To simulate this human review process, we execute tests to obtain feedback and identify unmet requirements, thereby evaluating the LLM’s IF capability in dynamic conversations. Specifically, we divide the entire session is into two types of rounds: instruction rounds and feedback rounds. In each instruction round, instructions that did not meet the requirements in the previous round are selected through testing. If the instruction round fails, it enters the feedback round for modification. To prevent the LLM from getting stuck in a loop, when both the instruction round and the corresponding feedback round fail, the instruction is discarded. Specifically, if $A_{i}$ does not meet the desired requirements, it undergoes a feedback round, where $I_{i+1}$ is constructed by incorporating $I_{i}$ and the feedback $F_{i}$ from the previous round. If $A_{i}$ meets the desired requirements, or if it cannot meet the corresponding desired based on feedback, then it enters the instruction round: by executing tests, select the instructions that the $A_{i}$ has not met.
Table 1. Overview of the verifiable instruction strategies of CodeIF-Bench. Two context-related strategies marked with * are customized for repository-level coding task.
@@
Code Complexity Control the complexity of the code structure (e.g., Cyclomatic Complexity (Ebert et al., 2016))
Code Standard Rules and guidelines for coding and structure (e.g., PEP 8 (Van Rossum et al., 2001))
@@
3.2. Evaluation Metrics
@@
  • Instruction Accuracy (IA) (He et al., 2024): This metric measures the percentage of instructions that the LLM accurately follows in each round of conversation. Specifically, in the $n$-th round of dialogue, given the historical dialogue and the current instruction $I_{n}$, the LLM outputs $A_{n}$. Whether the LLM complies with instruction $I_{n}$ is determined by judging whether $A_{n}$ passes the test $T_{n}$.
    (1) $IA=\begin{cases}1,&\text{if }A_{n}\text{ passes the test }T_{n}\\ 0,&\text{if }A_{n}\text{ does not pass the test }T_{n}\end{cases}$
@@
    (2) $CA=\frac{\text{The Number of Tests Passed by }A_{n}\text{ in }TS}{\text{The Number of Tests in TS}}$
  • Instruction Forgetting Ratio (IFR) (He et al., 2024): This metric assesses the percentage of instructions that are followed previously and not followed subsequently. Specifically, an instruction is considered forgotten if it was followed in one of the previous turns ($1,2,\dots,n-1$) but is not followed in the current turn ($n$). The test sequence passed in the previous turn $PTS=\{T^{\prime}_{1},...,T^{\prime}_{k}\}$ and the LLM outputs $A_{n}$.
    (3) $IFR=\frac{\text{The Number of Tests Failed by }A_{n}\text{ in }PTS}{\text{The Number of Tests in }PTS}$
  • Instructions Following Efficiency (IFE): This metric evaluates the instruction following efficiency of LLM in Dynamic Conversation. It represents the average number of instructions followed per round from the initial round to the current round in the current round. Specifically, in the $n$-th round of dialogue, given the historical dialogue $\{I_{1},A_{1},...I_{m},A_{m}\}$ and the current instruction $I_{n}$, the LLM outputs $A_{n}$. We calculate the IFE score in this turn by executing the test sequence $TS=\{T_{1},...,T_{m},T_{n}\}$.
    (4) $IFE=\frac{\text{The Number of Tests Passed by }A_{n}\text{ in }TS}{n}$
3.3. Benchmark Construction

### v4

@@
Abstract
Large Language Models (LLMs) have demonstrated exceptional performance in code generation tasks and have become indispensable programming assistants for developers. However, existing code generation benchmarks primarily assess the functional correctness of code generated by LLMs in single-turn interactions. They offer limited insight into LLMs’ abilities to generate code that strictly follows users’ instructions in multi-turn interaction scenarios. In this paper, we introduce CodeIF-Bench, a benchmark for evaluating the instruction-following capabilities of LLMs in interactive code generation. Specifically, CodeIF-Bench incorporates nine types of verifiable instructions aligned with the real-world software development requirements, which can be independently and objectively validated through specified test cases, facilitating the evaluation of instruction-following capability in multi-turn interactions. In both Static Conversation and Dynamic Conversation settings, we evaluate the performance of 6 state-of-the-art LLMs and summarize the important factors, additional repository context and gradually increasing interaction history influencing the instruction-following ability of LLMs in multi-turn interactions. Furthermore, we identify the potential direction for improvement: context management. The code and data are available at https://github.com/zhu-zhu-ding/CodeIF-Bench.
keywords:
@@ Abstract
1 Introduction
In recent years, the remarkable advancements in Large Language Models (LLMs) for code generation [zhu2024deepseek, roziere2023code, luo2023wizardcoder] have significantly enhanced developers’ coding efficiency through an interactive and conversational paradigm [GitHub-Copilot, Cursor].
@@ 1 Introduction
Figure 1: An example of interactive code generation, where the developer provides follow-up instructions to clarify the requirement and address issues in the generated code
In practical software development environments, as shown in Figure 1, the collaboration between developers and LLMs typically necessitates an interactive code generation process. Throughout this interactive workflow, developers frequently enhance and refine their initial instructions by adding supplementary details to better align with their requirements. Specifically, through systematic code execution and rigorous review processes, developers often identify and provide follow-up instructions to clarify requirements and address issues in LLM-generated code. These issues may include implementing new features [jimenez2023swe-bench], fixing bugs [codebugs], resolving potential security vulnerabilities [perry2023users], etc. Specifically, any discrepancies or misunderstandings in LLM following instructions can significantly increase interaction time and cost, as developers may need to spend additional time clarifying and fixing errors. Besides, such misunderstandings might even trigger uncontrollable and unexpected issues, such as bugs [codebugs], system failures [codeweaknesses], or even security vulnerabilities [perry2023users], which can have severe implications for the software’s performance and developers’ trust. Consequently, a dedicated benchmark to evaluate LLMs’ Instruction-Following (IF) capabilities in interactive code generation is essential.
The key challenges in benchmarking instruction following for interactive code generation lie in capturing the iterative, linguistic, and practical complexities of real-world developer–LLM collaboration:
  • Challenge 1 - Modeling multi-turn interactions. Unlike single-turn tasks, interactive code generation requires the benchmark to simulate iterative developer–LLM dialogues where each instruction depends on the previous interaction history. Constructing such multi-turn data is non-trivial, as it requires coherent context linking, and realistic instruction evolution.
  • Challenge 2 — Ambiguity and context dependence of instructions. Developers’ instructions are often ambiguous, partial, or context-dependent. Accurately determining whether an LLM has truly “followed” such instructions is challenging.
  • Challenge 3 — Ensuring real-world representativeness. To accurately assess LLM’s performance in real-world settings, the benchmark must reflect how real developers formulate and refine coding intents, rather than relying on generic natural language instructions that lack authentic development context and have already been extensively explored.
@@
For Challenge 1, existing popular code generation benchmarks [HumanEval, MBPP, APPS, DevEval] primarily evaluate whether LLMs can generate functionally correct code given specific programming tasks in a single turn, and it is difficult to directly evaluate LLMs’ instruction-following capability in multi-round interaction. For Challenge 2, for instance, subjective instructions such as “Please make the comments clearer” may result in varying interpretations when assessing whether the model follows the given requirement. Some instruction-following benchmarks in the NLP domain, including for multi-turn [multiif, MT-Bench, mt-bench-101], most use the LLM-as-Judge method for evaluation, which may lead to bias [li2025preferenceleakagecontaminationproblem, chen-etal-2024-humans]. To circumvent this challenge, some work, such as [IF-Bench], focuses on evaluating the ability of LLMs to follow “verifiable instructions” which are defined as instructions that can be objectively verified for compliance, e.g., “the response should be more than 300 words”. However, it does not address the Challenge 3: these benchmarks do not align with real-world software development scenarios and fail to evaluate LLM’s ability to follow instructions in realistic interactive coding tasks.
@@
To address the above challenges, we propose a new benchmark named CodeIF-Bench, which aims to evaluate LLMs’ instruction-following capability in interactive code generation. We first gather initial instructions from real-world projects [DevEval] and crowd-sourced problems [MBPP], enabling evaluation for both standalone function and repository-level code generation. Then we use a combination of LLM and manual annotation to extract nine primary verifiable instructions strategies of real-world user requirements to guide LLM to generate verifiable instructions, which can be independently and objectively verified using their respective test cases, such as initial instructions. We guide LLM to generate verifiable instructions based on these strategies and initial instructions, resulting in each data point in CodeIF-Bench containing at least 8 verifiable instructions (8 for function-level and 10 for repository-level) to construct a conversation. In addition, we design Static Conversation (predefined instruction sequences) and Dynamic Conversation (providing feedback information), and evaluation metrics based on test cases to evaluate the performance of LLM in code interaction scenarios.
We evaluate the IF capabilities of 6 widely-used and advanced LLMs (i.e. GPT-4o [openai2024gpt], Claude-3.5-Sonnet [claude], DeepSeek-V3 [zhu2024deepseek], Qwen2.5-Coder-{7B, 14B, 32B} [qwencoder]) in interactive code generation tasks using CodeIF-Bench. Experimental results reveal that as additional repository contexts and dialogue histories are incorporated, the instruction-following performance of LLMs gradually declines. This degradation manifests as a weakened ability to follow the current instruction and an increasing tendency to forget previous instructions. In dynamic interactive settings with feedback, the growing contextual burden further hinders LLMs from effectively leveraging feedback to improve their outputs. Our case studies and analyses of various prompting strategies further suggest that effective context management constitutes a promising direction for enhancing LLM performance.
In summary, our contributions are as follows:
@@
  • We perform a thorough evaluation of 6 prominent LLMs on CodeIF-Bench, analyzing the key factors influencing their IF capabilities in interactive code generation.
2 Related Work
Code Generation Benchmarks. Recent years have seen remarkable progress in LLM-based code generation, with models such as DeepSeek-Coder [zhu2024deepseek], WizardCoder [luo2023wizardcoder], and Qwen-Coder [qwencoder] demonstrating exceptional performance. This progress has facilitated the development of interactive coding tools [Cursor, GitHub-Copilot] that enhance programmer productivity through effective human-LLM collaboration. This development has driven the development of benchmarks in code generation. These include benchmarks for evaluating LLMs’ performance on various code generation scenarios, including stand-alone function generation (e.g., HumanEval [HumanEval], MBPP [MBPP], APPS [APPS]), class-level code generation (e.g., ClassEval [ClassEval]), and repository-level code generation (e.g., DevEval [DevEval] and CoderEval [CoderEval], which leverage real-world open-source repositories). While these benchmarks primarily assess whether LLMs can generate functionally correct code in single-turn interactions, they are insufficient for evaluating LLMs’ interactive capabilities, particularly in multi-turn interaction scenarios.
Instruction Following Benchmarks. Current benchmarks for evaluating the instruction-following capabilities of LLMs [IF-Bench, MT-Eval, MT-Bench, FB-Bench] primarily focus on natural language QA tasks in the NLP domain. However, there is a significant disparity exists between code generation and QA tasks. Some of the existing benchmarks utilize LLM-as-judge to assess LLMs’ instruction following performance [Systematic, huang2024empirical, chen2024mllm, MT-Bench], which typically rely on elaborate prompts and can introduce inherent bias. CanItEdit [CanItEdit] is a code editing benchmark, but its manually crafted programs lack project-specific dependencies, misaligning with real-world development. Additionally, its single-round dialogue design cannot assess LLMs’ performance in complex, multi-turn interactions. CodeIF [codeif] benchmarks instruction-following in code generation but oversimplifies real-world complexity. Its single-turn setting also fails to assess LLMs’ multi-turn interaction capabilities.
3 CodeIF-Bench
@@
Figure 2: CodeIF-Bench construction pipeline. The top part illustrates the verifiable instruction strategy extraction process, and the bottom part presents the data collection procedure
3.1 Task Definition
@@
For CodeIF-Bench, every data contains multiple verifiable instructions (VI) accompanied by test cases. Each verifiable instruction can be independently validated through tests to determine whether the model’s output follows the instruction. By linking instructions to their corresponding tests, CodeIF-Bench evaluates the LLMs’ instruction-following capabilities throughout the interaction process by executing these tests. Specially, each data in CodeIF-Bench is denoted as $D_{i}=\{(I_{1},T_{1}),(I_{2},T_{2}),...(I_{k},T_{k})\}$, $i\in\{1,2,...,|D|\},k\in\{1,2,...,|I|\},$ where $|D|$ is the total number of data and $|I|$ is the number of the verifiable instructions with tests. $I_{k}$ represents the $k$-th verifiable instruction, and $T_{k}$ represents the corresponding tests, which are both provided by the original benchmark or manual annotation. Based on CodeIF-Bench, we define the following two classic interactive code generation scenarios:
Static Conversation: In this scenario, we simulate a linear progression of requirements where instruction sequences are predetermined. The LLM generates code for each instruction in sequence without iterative refinement, testing its ability to maintain consistency in multi-turn interactions without external feedback. In the N-Round conversation, we initialize the process at $N=1$ with the original programming instruction ($I_{1}$) from the benchmark. Subsequent instructions are then added sequentially from a predefined non-conflicting instruction sequence as $N$ increases. In particular, for non-standalone tasks, which contain context-aware dependencies in repositories, $I_{1}$ consists of both the original task instruction and the golden contexts. Specifically, in $i$-th round, the input is formed as $\textit{Input}=(I_{1},A_{1},I_{2},A_{2},...,I_{i})$, LLM’s answer in this round is $\textit{A}_{i}$ evaluated with the test cases $(T_{1},T_{2},...,T_{i})$.
Dynamic Conversation: In real-world coding scenario, programmers typically review LLM-generated outputs to provide feedback. To simulate this human review process, we execute tests to obtain feedback and identify unmet requirements, thereby evaluating the LLM’s IF capability in dynamic conversations. Specifically, each interaction round consists of both an instruction and its corresponding feedback, forming an instruction–feedback pair. We first initialize the instruction list containing all verifiable instructions and tests $(I_{1},T_{1},\dots,I_{n},T_{n})$. During $i$ round, the system first executes the generated answer $A_{i-1}$ according to the current instruction list with tests. If $A_{i-1}$ fails to meet the requirements $(I_{m},T_{m})$, the instruction in this round is the $I_{m}$ with feedback $F_{m}$ to get the answer $A_{i}$. If $A_{i}$ passes the tests $T_{m}$, we keep the $(I_{m},T_{m})$ in current instruction list. Otherwise, we remove the $(I_{m},T_{m})$ from the instruction list. Until the instruction list is empty, the maximum number of interaction rounds is reached or the answer passes all instructions.
Table 1: Overview of the verifiable instruction strategies of CodeIF-Bench. Two context-related strategies marked with * are customized for repository-level coding task
@@
Code Complexity Control the complexity of the code structure (e.g., Cyclomatic Complexity [Cyclomatic])
Code Standard Rules and guidelines for coding and structure (e.g., PEP 8 [pep8])
@@
3.2 Evaluation Metrics
In interactive code generation, we evaluate LLMs from three key perspectives. First, LLMs should not only follow the current instruction but also maintain consistent comprehension and execution across the entire dialogue. Second, we analyze the instruction forgetting phenomenon, where previously followed instructions are no longer respected in later turns. Third, in the Dynamic Conversation setting, where feedback is introduced, we assess instruction-following efficiency—i.e., how many interaction rounds are required to complete $N$ instructions. Ideally, this would take exactly $N$ rounds, but due to corrective feedback, it may require $2N$ or more.
  • Instruction Accuracy (IA) [multiif]: This metric measures the percentage of instructions that the LLM accurately follows in each round of conversation. Specifically, in the $n$-th round of dialogue, given the historical dialogue and the current instruction $I_{n}$, the LLM outputs $A_{n}$. Whether the LLM complies with instruction $I_{n}$ is determined by judging whether $A_{n}$ passes the test $T_{n}$.
@@ 3.2 Evaluation Metrics
[…]
  • Instruction Forgetting Ratio (IFR) [multiif]: This metric assesses the percentage of instructions that are followed previously and not followed subsequently. Specifically, an instruction is considered forgotten if it was followed in one of the previous turns ($1,2,\dots,n-1$) but is not followed in the current turn ($n$). The test sequence passed in the previous turn $PTS=\{T^{\prime}_{1},...,T^{\prime}_{k}\}$ and the LLM outputs $A_{n}$.
[…]
  • Cumulative Instruction Following (CIF): This metric evaluates the cumulative instruction following of LLM in Dynamic Conversation. It represents the number of instruction followed at the conclusion of the entire dynamic session. Specifically, in the last round of dialogue, given the test sequence $TS=\{T_{1},...,T_{n}\}$ in the data, the LLM outputs $A_{last}$. We calculate the CIF score in this turn by executing the test sequence.
[…]
3.3 Benchmark Construction